# GRAFT: Gradient-Aware Fast MaxVol Technique for Dynamic Data Sampling

### v1

@@ 3.3 Complexity Analysis
We establish convergence guarantees for GRAFT under standard assumptions: (A1) $\mathcal{L}$ is $L$-smooth; (A2) stochastic gradients are unbiased with bounded variance; (A3) projection error satisfies $\|\bar{g}_{i}-G_{R}G_{R}^{\dagger}\bar{g}_{i}\|_{2}\leq\varepsilon$.
@@
Under (A1)–(A3), SGD with projected directions $G_{R}G_{R}^{\dagger}\bar{g}_{i}$ converges to a stationary point with $\min_{t\leq T}\mathbb{E}\big{\|}\nabla\mathcal{L}(\Theta^{t})\big{\|}_{2}^{2}\leq O\!\left(\frac{1}{\sqrt{T}}\right)+O(\varepsilon^{2})$
Theorem 3 (Computational complexity).
Fast MaxVol selection with gradient-projection evaluation costs $O(Kr_{\max}d_{z})+O(KR^{2})+O(pR^{2})$, where $K$ is batch size, $R$ is selected rank, $d_{z}$ is feature dimension, and $p=|\Theta|$. Cost is independent of dataset size $N$.

### v2

@@ 3.3 Complexity Analysis
Under (A1)–(A3), SGD with projected directions $G_{R}G_{R}^{\dagger}\,\bar{g}_{i}$ satisfies
$\min_{t\leq T}\ \mathbb{E}\bigl{\|}\nabla L(\Theta_{t})\bigr{\|}_{2}^{2}\;\leq\;O\!\left(\tfrac{1}{\sqrt{T}}\right)\;+\;O(\varepsilon^{2}).$
Proof sketch. With $L$ smooth and stochastic gradients unbiased with bounded variance, the standard SGD descent recursion yields an $O(1/\sqrt{T})$ rate. The projection step introduces a deterministic bias whose energy is controlled by the bound $\|\bar{g}_{i}-G_{R}G_{R}^{\dagger}\bar{g}_{i}\|_{2}^{2}\leq\varepsilon^{2}$. This adds an $O(\varepsilon^{2})$ term to the stationarity gap while leaving the stochastic term unchanged; see the Supplement (Detailed Proofs) for a full derivation. $\square$
Computational complexity Per iteration, Fast MaxVol selection with gradient–projection evaluation runs in
$O\!\left(KR^{2}\right)\;+\;O\!\left(|\text{Rset}|\,R\,d\right)$
time and uses $O(Kd+dR+R^{2})$ memory. Computing gradients for the selected $R$ samples contributes $O(Rd)$ and is included when reported separately. A basis/feature refresh (performed periodically) costs $O(KdR)+O((K{+}d)R^{2})$ when it occurs and is amortized over the refresh interval. The per-iteration term is independent of $N$. Fast MaxVol scans $K$ rows of the $K{\times}R$ feature matrix while maintaining rank-1 volume/determinant updates, giving $O(KR^{2})$. For rank selection, evaluating a projection criterion (e.g., $\|\bar{g}-P_{r}\bar{g}\|_{2}$) over $r\in\text{Rset}$ costs $\sum_{r\in\text{Rset}}O(rd)=O(|\text{Rset}|\,R\,d)$. Gradients for the $R$ chosen samples cost $O(Rd)$ if accounted separately. The feature/basis refresh is performed only at scheduled steps and thus amortizes; all operations act on the current mini-batch, so no term scales with $N$. Full details are in the Supplement (Complexity Analysis). $\square$